Write 34/4 as a mixed fraction


1
Expert's answer
2021-05-07T09:22:02-0400

Solution. Divide 34 by 4


"34=4\times8 +2."

As result get


"\frac{34}{4}=8\frac{2}{4}=8\frac{1}{2}"

Answer.

"\frac{34}{4}=8\frac{1}{2}"